Test: Exam#2 ti This Question: 1 pt 5 of 19 (10 complete) Let p be the population proportion for the following condition. Find the point estimates for p and q. A study of 4490 adults from country A, 3024 think mainstream media is more interested in making money than in telling the truth. The point estimate for p, p. is Round to three decimal places as needed.) The point estimate for q. 4 is Round to three decimal places as needed.)
Answer #1

Here we have

x=3024 and n=4490

(a) Point estimate for p is p^=x/n=3024/4490= 0.673

(b) Point estimate for q is q^= 1-p=1-0.673=0.327